# PTEB: Towards Robust Text Embedding Evaluation via Stochastic Paraphrasing at Evaluation Time with LLMs

### v1

@@ 3.2.3 Embedding Model Evaluation on PTEB
$\Delta$-3.17/±2.23 $\Delta$-5.60/±1.61 $\Delta$-2.03/±1.11 $\Delta$-6.93/±1.01 $\Delta$-4.85/±1.53
@@
$\Delta$-3.15/±0.09 $\Delta$-1.71/±0.07 $\Delta$-4.55/±0.07 $\Delta$-1.75/±0.05 $\Delta$-3.09/±0.07
@@
$\Delta$-10.79/±0.29 $\Delta$-8.58/±0.41 $\Delta$-11.16/±0.24 $\Delta$-7.60/±0.32 $\Delta$-14.20/±0.30
@@
$\Delta$-4.58/±0.56 $\Delta$+5.99/±0.64 $\Delta$-5.62/±0.33 $\Delta$+0.04/±0.35 $\Delta$-4.99/±0.41
@@
$\Delta$-5.86/±0.26 $\Delta$+0.08/±0.19 $\Delta$-6.81/±0.12 $\Delta$-1.14/±0.21 $\Delta$-6.22/±0.22
@@
$\Delta$-4.08/±0.11 $\Delta$-2.72/±0.19 $\Delta$-4.29/±0.20 $\Delta$+1.00/±0.10 $\Delta$-3.06/±0.12
@@
$\Delta$-5.55/±0.25 $\Delta$-7.33/±0.70 $\Delta$-4.22/±0.28 $\Delta$-3.25/±0.31 $\Delta$-3.59/±0.21
@@
$\Delta$+1.08/±0.48 $\Delta$+5.16/±0.67 $\Delta$-0.05/±0.38 $\Delta$-2.17/±0.74 $\Delta$-1.37/±0.66
@@
$\Delta$-8.53/±0.35 $\Delta$-0.48/±0.53 $\Delta$-7.09/±0.36 $\Delta$-2.56/±0.38 $\Delta$-3.94/±0.32
@@
$\Delta$-4.96 $\Delta$-1.69 $\Delta$-5.09 $\Delta$-2.70 $\Delta$-5.03
Table 5: Embedding model performance on original and PTEB STS datasets (in %). Each cell shows the original and PTEB score; the difference is denoted by $\Delta$. Red indicates performance decreases; green improvements. All values are aggregated over $n=6$ runs with the sample standard deviation over the different paraphrases denoted by ±. Bold marks row-wise best for original and paraphrased text separately.
To test statistical significance, we use the paired Wilcoxon signed-rank test, which can be applied in many NLP settings (Dror et al., 2018). When pooling over all STS datasets, all embedding models except for embeddinggemma-300m outperform the baseline all-mpnet-base-v2 significantly after Holm-correction (Holm, 1979) ($p\ll\alpha=0.05$; for all models: $0\notin\text{CI}_{95\%}(\Delta_{HL})$; see Table 6). [^3] The Holm-method adjusts p-values when testing multiple hypotheses. It sorts the $m$ hypotheses by their p-values ($p_{i}$) and then checks, starting with the lowest p-value, for a given significance level $\alpha$ if $\forall i\in\{1,\dots,m\}:p_{i}<\frac{\alpha}{m-i+1}$ (Calonico and Galiani, 2025). The estimated location shifts $\Delta_{HL}$ [^4] The Hodges–Lehmann estimator (Hodges and Lehmann, 1963) associated with the Wilcoxon signed-rank test is the median of all Walsh averages, i.e., the pairwise means formed from every self-pair and unique unordered pair of the observed within-pair differences. are negative for all models except embeddinggemma-300m.
mpnet vs. $\Delta_{HL}$ [CI] p-value
embedding- +7.33 6.50e-10
gemma-300m [7.01, 8.12]
mxbai-embed- -3.74 1.33e-08
large-v1 [-5.11, -3.09]
e5-mistral- -3.48 1.40e-07
7b-instruct [-4.37, -2.60]
qwen3-em- -4.53 1.48e-09
bedding-8b [-4.82, -4.19]
Table 6: Comparisons on PTEB STS vs. all-mpnet-base-v2. $\Delta_{HL}$ is the Hodges–Lehmann estimator (Hodges and Lehmann, 1963) with its 95% confidence interval; negative values favour the listed model. P-values are Holm-adjusted. ($n=6$ runs)
3.2.4 Ablation on STS Embedding Model Size
To analyse the influence of the encoder model size, we conduct an ablation study using the Qwen3 models (Table 7). The performance gap between the original and paraphrased datasets remains similar across all three Qwen3 variants, with the smallest model even displaying the narrowest gap.
STS Score Embedding Model Size
@@ 3.2.4 Ablation on STS Embedding Model Size
Original 81.42 84.26 84.90
PTEB 77.19 78.72 79.87
@@
Task
all-mpnet-
base-v2
embedding
gemma-300m
mxbai-embed-
large-v1
e5-mistral-
7b-instruct
qwen3-em-
bedding-8b
@@
Classification 92.23 85.21 85.83 80.56 93.66 87.42 90.99 84.56 92.78 86.64
Banking77 $\Delta$ -7.02 $\Delta$ -5.27 $\Delta$ -6.24 $\Delta$ -6.43 $\Delta$ -6.14
Clustering 50.18 48.24 32.99 34.33 51.25 48.93 45.61 39.04 55.51 49.46
TwentyNewsGroups $\Delta$ -1.94 $\Delta$ +1.34 $\Delta$ -2.32 $\Delta$ -6.57 $\Delta$ -6.05
Pair Classif. 73.85 71.36 59.03 63.73 78.54 73.11 75.36 73.78 72.22 73.90
TwitterSemEval $\Delta$ -2.49 $\Delta$ +4.70 $\Delta$ -5.43 $\Delta$ -1.58 $\Delta$ +1.68
Reranking 65.82 64.42 48.14 48.20 65.20 63.98 59.96 59.51 67.48 65.27
AskUbuntuDupQ. $\Delta$ -1.40 $\Delta$ +0.06 $\Delta$ -1.22 $\Delta$ -0.45 $\Delta$ -2.21
Retrieval 45.89 45.45 29.66 32.89 65.11 61.18 53.52 49.89 75.27 73.62
ArguAna $\Delta$ -0.44 $\Delta$ +3.23 $\Delta$ -3.93 $\Delta$ -3.63 $\Delta$ -1.65
Summarisation 26.25 23.74 20.35 15.74 35.55 32.59 34.24 32.80 38.87 35.29
SummEval $\Delta$ -2.51 $\Delta$ -4.61 $\Delta$ -2.96 $\Delta$ -1.44 $\Delta$ -3.58
Average 59.04 56.40 46.00 45.91 64.88 61.20 59.95 56.60 67.02 64.03
$\Delta$ -2.64 $\Delta$ -0.09 $\Delta$ -3.68 $\Delta$ -3.35 $\Delta$ -2.99
Table 8: Embedding model performance for non-STS tasks and datasets on original data and PTEB. Each cell shows the original and PTEB score; the difference is denoted by $\Delta$. Performance decreases red; improvements green. Bold: best per column. See Appendix A for metrics per dataset and references. ($n=1$ runs; in %)
Model Classification Pair Class. STS
Original PTEB Original PTEB Original PTEB
paraphrase-multilingual-mpnet-base-v2 88.54 85.55 89.14 89.29 83.46 79.98
embeddinggemma-300m 90.51 86.11 86.95 86.49 61.89 59.74
multilingual-e5-large-instruct 89.08 86.32 87.62 87.23 72.82 70.86
qwen3-embedding-8b 89.83 87.04 87.35 87.24 83.35 80.24
Avg. word count text 1/text 2 19.9 18.8 30.5/8.2 30.3/9.4 8.3/8.1 8.8/8.4
Table 9: Multilingual eval: AmazonCounterFactuals (class.), RTE3 (pair class.), STS17 (STS) ($n=6$ runs; in %). Performance drops red; improvements green. Bold: best per column. For paired data, we report both word counts.
Paraphrase Prompt
PTEB STS
Score
Edit
Distance
Default 74.92 80.87
Variant 1 74.24 81.49
Variant 2 74.82 82.00
Variant 3 75.16 83.63
Table 10: PTEB STS scores and the corresponding edit distances to the original text for the default paraphrase prompt and 3 differently phrased prompt variants; see Appendix C.3 for the prompts used. (Based on all-mpnet-base-v2 embeddings; $n=1$ runs; in %)
@@ 3.2.5 PTEB for Non-STS Tasks
Extending PTEB to the remaining 6 MTEB tasks, shows that the average performance of all models drops on PTEB compared to original texts (Table 8). qwen3-8b is the best-performing model on 5 out of the 6 tasks. The total performance drop of embeddinggemma-300m is marginal (-0.09) while the other models drop between 2.64 and 3.35. (For an additional visualisation see Appendix Figure 6.)
@@ 3.2.6 Multilingual PTEB
We validate our results on non-English data, using RTE3 (Giampiccolo et al., 2007), AmazonCounterFactuals (O’Neill et al., 2021), and the multilingual version of STS17; covering 10 languages (see Table 9 for results; see footnote 6 for a list of language pairs). We used the same models as before if they are natively multilingual, or their multilingual variants alternatively. We find significant performance drops on AmazonCounterFactuals and STS17 while the results are stable on RTE3.
@@ 3.2.7 STS Paraphrase Prompting Sensitivity
Since LLMs can be sensitive to prompt changes (Mizrahi et al., 2024), we tested the influence of 3 variations of the paraphrase prompt and compared the PTEB STS scores and the ED between original and paraphrased texts (Table 10). We find limited differences ($\leq 0.68\%$ for the score and $\leq 2.76\%$ for the ED).
@@ 4 Discussion
Our method demonstrates that the 3 generative LLMs evaluated in our study are well-suited to rate semantic similarity with STS scores above 80% (Table 2). Gemma-3-27b even outperformed 3 of the 5 embedding models on average on the original STS data (see Table 2 vs. Table 5). [^5] With the limitation that Gemma-3-27b has been evaluated only on one run and is hence not a like-for-like comparison. Furthermore, the evaluated LLMs generate paraphrases with high semantic similarity while introducing token-level variation from the original text (Table 4). We found that gemma-3-27b generated more textually diverse paraphrases at higher throughput: average ED 0.81 vs. 0.55 and 0.63 at a 5× to 20× speed-up compared to gpt-oss-20b and qwen3-32b respectively.
Moreover, across the 7 PTEB tasks and 10 languages, most of the 18 datasets show decreased performance of embedding models on paraphrased data (Table 5, Table 8, Table 9). Our results support the hypothesis of some benchmark saturation and overfitting. While qwen3-8b achieves the highest overall scores on the original data and PTEB, reordering occurs at the dataset level. qwen3-8b, for example, performs best on 4 PTEB datasets where it does not achieve the highest score on the original data, and vice versa; it ranks highest on 3 original datasets where it does not lead on PTEB. This demonstrates that embedding models show different degrees of robustness towards paraphrasing.
For STS22, the relatively small performance drops on PTEB may be explained by its higher word count (avg. 461 per text vs. 61 for all STS data; see Table 1). Based on qualitative observations, we suggest that STS22’s longer texts may allow paraphrases with fewer lexical changes, while paraphrases of shorter sentences rely more on synonym replacement. Moreover, on shorter sentences single tokens have more weight and may distort the embedding model. Another outlier is RTE3 with only marginal score changes. This might stem from its intuitively more robust entailment relation between a long text 1 and a short text 2 (Table 9).
@@
Furthermore, English datasets with initial scores ($\geq 80$), such as BIOSSES, SICK-R, STS13/15/17/B, and Banking77 showed relatively large performance gaps (many models -3 to -6%; also see Appendix Figure 6 which visualises this on task level). This suggests that the initially high scores may depend on shortcuts in token space which are neutralised by paraphrasing. It shows that PTEB’s new evaluation method adds a stress test for semantic invariance, mitigating concerns that MTEB may favour lexical shortcuts and overfitted embedding models. However, neither do we claim that any specific model is more overfitted than another one nor do we make general claims on the performance of individual models as the aim of our study is to present a novel evaluation method; not a comprehensive benchmarking study.
@@
Unexpectedly, smaller encoders were not more sensitive to paraphrasing: Our ablation study on encoder size showed the similar performance and robustness of qwen-3-4b vs. the 8b model with reduced runtime and VRAM (Table 7). For the datasets in our study, the results indicate that robust performance under paraphrase stress can be achieved with smaller embedding models. Furthermore, the prompt sensitivity analysis showed that PTEB is robust to prompt changes further mitigating potential concerns that the results might be brittle due to the use of generative LLMs (Table 10).

### v2

@@ 3.2.3 PTEB for STS
$\Delta$-3.17 / ±2.23 $\Delta$-5.60 / ±1.61 $\Delta$-2.03 / ±1.11 $\Delta$-6.93 / ±1.01 $\Delta$-4.85 / ±1.53 $\Delta$-4.52
@@
$\Delta$-3.15 / ±0.09 $\Delta$-1.71 / ±0.07 $\Delta$-4.55 / ±0.07 $\Delta$-1.75 / ±0.05 $\Delta$-3.09 / ±0.07 $\Delta$-2.85
@@
$\Delta$-10.79 / ±0.29 $\Delta$-8.58 / ±0.41 $\Delta$-11.16 / ±0.24 $\Delta$-7.60 / ±0.32 $\Delta$-14.20 / ±0.30 $\Delta$-10.47
@@
$\Delta$-4.58 / ±0.56 $\Delta$+5.99 / ±0.64 $\Delta$-5.62 / ±0.33 $\Delta$+0.04 / ±0.35 $\Delta$-4.99 / ±0.41 $\Delta$-1.83
@@
$\Delta$-5.86 / ±0.26 $\Delta$+0.08 / ±0.19 $\Delta$-6.81 / ±0.12 $\Delta$-1.14 / ±0.21 $\Delta$-6.22 / ±0.22 $\Delta$-3.99
@@
$\Delta$-4.08 / ±0.11 $\Delta$-2.72 / ±0.19 $\Delta$-4.29 / ±0.20 $\Delta$+1.00 / ±0.10 $\Delta$-3.06 / ±0.12 $\Delta$-2.63
@@
$\Delta$-5.55 / ±0.25 $\Delta$-7.33 / ±0.70 $\Delta$-4.22 / ±0.28 $\Delta$-3.25 / ±0.31 $\Delta$-3.59 / ±0.21 $\Delta$-4.79
@@
$\Delta$+1.08 / ±0.48 $\Delta$+5.16 / ±0.67 $\Delta$-0.05 / ±0.38 $\Delta$-2.17 / ±0.74 $\Delta$-1.37 / ±0.66 $\Delta$+0.53
@@
$\Delta$-8.53 / ±0.35 $\Delta$-0.48 / ±0.53 $\Delta$-7.09 / ±0.36 $\Delta$-2.56 / ±0.38 $\Delta$-3.94 / ±0.32 $\Delta$-4.52
@@
$\Delta$-4.96 $\Delta$-1.69 $\Delta$-5.09 $\Delta$-2.70 $\Delta$-5.03 $\Delta$-3.90
Table 6: Embedding model performance on original and PTEB STS datasets. $\Delta$ denotes the difference with Red indicating performance decreases and Green improvements. Sample standard deviation over the different paraphrases denoted by ±. Bold marks row-wise best for original and paraphrased text separately. ($n=6$, in %)
all-mpnet-base-v2 vs. $\Delta_{HL}$ [CI] $p_{Holm}$
embedding-gemma-300m +7.33 6.50e-10
[7.02, 8.00]
mxbai-embed-large-v1 -3.74 1.33e-08
[-5.00, -3.14]
e5-mistral-7b-instruct -3.48 1.40e-07
[-4.37, -2.56]
qwen3-embedding-8b -4.53 1.48e-09
[-4.82, -4.19]
Table 7: Comparison of encoders on PTEB STS vs. all-mpnet-base-v2 as baseline. $\Delta_{HL}$ is the Hodges-Lehmann estimator of the location shift with 95% CI; negative values favour the listed model. P-values are Holm-adjusted. Based on $9\times 6$ dataset-run pairs.
On average, all models decline on PTEB vs. the original STS datasets with the smallest drop for embeddinggemma-300m (Table 6). e5-mistral-7b-instruct decreases by 2.70%, while others drop by up to 5.09%. The gap is largest on STS12 (-10.47) while STS22 shows a small average improvement of +0.53.
Embedding Model Size
@@
Original STS Score 81.42 84.26 84.90
PTEB STS Score 77.19 78.72 79.87
@@
Task all-mpnet- base-v2 embedding gemma-300m mxbai-embed- large-v1 e5-mistral- 7b-instruct qwen3-em- bedding-8b Average
@@
STS 9 datasets 80.35 75.39 69.31 67.62 84.43 79.34 81.45 78.75 84.90 79.87
(see Table 6) $\Delta$-4.96 / ±1.37 $\Delta$-1.69 / ±0.49 $\Delta$-5.09 / ±0.40 $\Delta$-2.70 / ±0.43 $\Delta$-5.03 / ±0.42 $\Delta$-3.89
Classification 92.23 85.90 85.83 81.16 93.66 87.86 90.99 84.90 92.85 86.77
Banking77 $\Delta$ -6.33 / ±0.48 $\Delta$ -4.67 / ±0.37 $\Delta$ -5.80 / ±0.39 $\Delta$ -6.09 / ±0.43 $\Delta$ -6.08 / ±0.37 $\Delta$-5.79
Clustering 50.18 47.44 32.99 34.51 51.25 48.52 45.61 39.57 55.51 50.03
TwentyNewsGr. $\Delta$ -2.74 / ±0.71 $\Delta$ +1.52 / ±0.51 $\Delta$ -2.73 / ±0.56 $\Delta$ -6.04 / ±0.99 $\Delta$ -5.48 / ±0.75 $\Delta$-3.09
Pair Classif. 73.85 71.65 59.03 63.90 78.54 73.65 75.36 74.40 72.22 74.21
TwitterSemEval $\Delta$ -2.20 / ±0.39 $\Delta$ +4.87 / ±0.10 $\Delta$ -4.89 / ±0.37 $\Delta$ -0.96 / ±0.49 $\Delta$ +1.99 / ±0.45 $\Delta$-0.64
Reranking 65.81 64.30 48.53 48.86 65.20 63.78 59.96 59.31 67.55 64.98
AskUbuntuDupQ. $\Delta$ -1.51 / ±0.40 $\Delta$ +0.33 / ±0.53 $\Delta$ -1.42 / ±0.30 $\Delta$ -0.65 / ±0.46 $\Delta$ -2.57 / ±0.26 $\Delta$-1.16
Retrieval 45.82 45.38 29.65 33.05 65.11 61.36 53.29 49.54 75.26 73.30
ArguAna $\Delta$ -0.44 / ±0.28 $\Delta$ +3.40 / ±0.28 $\Delta$ -3.75 / ±0.20 $\Delta$ -3.75 / ±0.38 $\Delta$ -1.96 / ±0.23 $\Delta$-1.30
Summarisation 26.25 23.55 20.35 15.57 35.55 33.02 34.24 32.27 38.87 32.95
SummEval $\Delta$ -2.70 / ±0.65 $\Delta$ -4.78 / ±0.58 $\Delta$ -2.53 / ±0.46 $\Delta$ -1.97 / ±1.15 $\Delta$ -5.92 / ±1.27 $\Delta$-3.58
Average 62.07 59.09 49.38 49.24 67.68 63.93 62.99 59.82 69.59 66.02
$\Delta$ -2.98 $\Delta$ -0.15 $\Delta$ -3.74 $\Delta$ -3.17 $\Delta$ -3.58 $\Delta$-2.72
Table 9: Embedding model performance across STS and Non-STS tasks (English); the STS total (from Table 6) is included as the first row. $\Delta$ denotes differences, Red marks decreases and green improvements. ± denotes the standard deviation on PTEB. Bold marks row-wise best for Original and PTEB separately. ($n=6$ runs; in %)
To test statistical significance, we use the Wilcoxon signed-rank test, which is widely applicable in NLP (Dror et al., 2018). [^5] Following concerns that bootstrapping requires larger sample sizes (Søgaard et al., 2014; Hesterberg, 2015), we avoid it for significance testing. Pooling over the $9\times 6$ STS dataset-run pairs, the negative HL location shifts ($\Delta_{HL}$) and Holm-corrected (Holm, 1979) [^6] The Holm-method adjusts p-values when testing multiple hypotheses. It sorts the $m$ hypotheses by their p-values ($p_{i}$) and then checks, starting with the lowest p-value, for a given significance level $\alpha$ if $\forall i\in\{1,\dots,m\}:p_{i}<\frac{\alpha}{m-i+1}$ (Calonico and Galiani, 2025)). p-values $\ll\alpha=0.05$ show that all embedding models except embeddinggemma-300m outperform the baseline significantly (see Table 7; for a more conservative analysis averaging across runs before testing, see Table 14 in Appendix E).
3.2.4 Ablation on STS Embedding Model Size
To analyse the influence of the size of the encoder, we conducted an ablation study using the Qwen3 models (Table 8). The performance gap between the original and paraphrased datasets remains similar across all 3 Qwen3 variants, with the smallest model even displaying the narrowest gap.
@@ 3.2.5 PTEB for Non-STS Tasks
We extend PTEB to 6 additional MTEB tasks in order to incorporate a broad range of tasks ranging from straightforward binary decisions to nuanced clustering of sentences. Table 9 shows that the average performance of all models drops on PTEB compared to original texts, although particularly embeddinggemma-300m improves on multiple tasks. qwen3-8b is the best-performing model in 5 of the 7 tasks on original data and 6 of 7 on PTEB. (For an additional visualisation see Appendix F.)
Task paraphrase-multi- lingual-mpnet-base-v2 embedding gemma-300m multilingual-e5- large-instruct qwen3-em- bedding-8b Average
Orig. PTEB Orig. PTEB Orig. PTEB Orig. PTEB
Classification 88.54 85.55 90.51 86.11 89.08 86.32 89.83 87.04
AmazonCounterf. $\Delta$ -2.99 / ±0.42 $\Delta$ -4.40 / ±0.36 $\Delta$ -2.77 / ±0.17 $\Delta$ -2.79 / ±0.31 $\Delta$-3.24
Clustering 18.98 19.41 16.17 15.51 24.77 21.04 25.49 24.42
MasakhaNEWS $\Delta$ +0.43 / ±0.87 $\Delta$ -0.67 / ±1.43 $\Delta$ -3.72 / ±2.31 $\Delta$ -1.08 / ±1.53 $\Delta$-1.26
Pair Classifi- 89.14 89.29 86.95 86.49 87.62 87.23 87.35 87.24
cation RTE3 $\Delta$ +0.15 / ±0.05 $\Delta$ -0.46 / ±0.24 $\Delta$ -0.39 / ±0.13 $\Delta$ -0.11 / ±0.18 $\Delta$-0.20
Reranking 58.74 58.93 22.62 24.11 69.51 69.80 75.07 75.08
RuBQReranking $\Delta$ +0.19 / ±0.68 $\Delta$ +1.49 / ±0.60 $\Delta$ +0.29 / ±0.13 $\Delta$ +0.01 / ±0.30 $\Delta$+0.50
Retrieval 58.26 57.78 29.15 25.70 72.62 73.16 78.72 80.54
TwitterHjerne $\Delta$ -0.48 / ±1.21 $\Delta$ -3.45 / ±0.63 $\Delta$ +0.54 / ±1.18 $\Delta$ +1.82 / ±0.89 $\Delta$-0.40
STS 83.46 79.98 61.89 59.74 72.82 70.86 83.35 80.24
STS17 $\Delta$ -3.48 / ±0.19 $\Delta$ -2.15 / ±0.51 $\Delta$ -1.96 / ±0.26 $\Delta$ -3.11 / ±0.20 $\Delta$-2.67
Average 66.19 65.16 51.22 49.61 69.40 68.07 73.30 72.43
$\Delta$ -1.03 $\Delta$ -1.61 $\Delta$ -1.33 $\Delta$ -0.88 $\Delta$-1.21
Table 10: Multilingial embedding model performance on original and PTEB datasets (25 languages); $\Delta$ denotes differences, Red marks decreases and green improvements. ± denotes the standard deviation on PTEB. Bold marks row-wise best for Original and PTEB separately. ($n=6$ runs; in %)
@@ 3.2.6 Multilingual PTEB
We validate our results on 6 datasets from 6 different tasks that include non-English data (incl. the multilingual version of STS17) covering 25 languages. We examine both widely-studied languages (e.g. Arabic) and under-resourced ones (e.g. Danish) incl. 13 from sub-Saharan Africa (e.g. Suwaheli, Hausa, Amharic) to evaluate performance across linguistic boundaries (see Table 10 for results; see Appendix A for the languages included). [^7] We did not include a non-English dataset for the summarisation task, as MTEB does not provide one. We used the same encoders as before if they are natively multilingual, or their multilingual variants alternatively. Consistent with English datasets, we find the largest drops on the classification task followed by STS, and clustering. An exception is the Russian RubQ dataset on which all models improve. As expected, for each task the average decreases are more pronounced on the English datasets compared to non-English datasets.
3.2.7 Human Evaluation
Languages for human evaluation were selected to (1) cover high and low resource languages (including low-resource Sub-Saharan languages like Hausa and Swahili), (2) represent a wide range of PTEB tasks (spanning Classification, Clustering, Pair Classification, Reranking, and STS), and (3) encompass diverse language families (e.g., Afroasiatic, Indo-European, Niger-Congo, and Turkic).
Language No. of Ratings Mean Similarity Median Similarity
Arabic (ARA) 2x50 4.43 5.00
French (FRA) 2x50 3.78 4.00
German (DEU) 2x50 4.40 5.00
Hausa (HAU) 2x50 3.34 4.00
Russian (RUS) 2x50 4.16 5.00
Spanish (SPA) 2x50 4.01 4.00
Swahili (SWA) 2x50 3.91 4.00
Turkish (TUR) 2x50 4.40 5.00
Total 2x400 4.05 4.00
Table 11: Human evaluation of the similarity between original texts and paraphrases (on scale from 0 to 5).
The human similarity ratings ranged from 3.34 (Hausa) to 4.43 (Arabic), with an average of 4.05 and a pooled median of 4.00 (Table 11). This confirms a high degree of semantic similarity between the generated paraphrases and the original text. All languages exceeded the mean acceptability threshold. While the majority of sentence pairs were rated 3.00 or higher, 13% of sentence pairs in Hausa received a similarity score of 0.00 (Figure 6).
Figure 6: Human paraphrase similarity rating distribution (based on $n=2\times 50$ ratings per language).
@@ 3.2.8 STS Paraphrase Prompting Sensitivity
Since LLMs can be sensitive to prompt changes (Mizrahi et al., 2024), we tested the influence of 3 variations of the paraphrase prompt and compared the PTEB STS scores and the ED between original and paraphrased texts (Table 12). We find limited differences ($\leq 0.92\%$ for the score and $\leq 0.03$ for the ED).
Paraphrase Prompt PTEB STS Score Edit Distance
Default 74.92 0.81
Variant 1 74.24 0.82
Variant 2 74.82 0.82
Variant 3 75.16 0.84
Table 12: PTEB STS scores (English; in %) and the corresponding edit distances to the original text for the default paraphrase prompt and 3 differently phrased prompt variants; see Appendix D.4 for the prompts. (Based on all-mpnet-base-v2 embeddings; $n=1$ runs)
@@ 4 Discussion
[…]
Based on our human evaluation for 8 languages, we found that the generated paraphrases are highly similar to the original sentences (mean: 4.05; median: 4.00; see Table 11). However, for Hausa, Swahili, and French, only 70%, 84%, and 84% of sentence pairs met the similarity threshold of 3.00. While this may indicate limited paraphrasing capabilities for low-resource languages like Hausa and Swahili, it is surprising that a high-resource language like French exhibits similar issues.
Across the 25 languages, we observe that all models exhibit lower average scores on PTEB. This drop is most pronounced in English (-2.72pt vs. -1.21pt; Table 9 vs. Table 10), suggesting that encoders (particularly high-performing ones) may be overfitted to MTEB, relying on shortcuts in the token space rather than semantic understanding. Classification tasks show the largest average performance declines across both English and non-English datasets. We hypothesise that this occurs because classification relies heavily on specific keyword triggers that are disrupted by paraphrasing.
Except for STS22 and RubQ, 18 out of the 20 datasets show lower PTEB scores averaged over all embedding models vs. original scores. For STS22, this can largely be attributed to Embedding-Gemma while the relatively small performance drops for other encoders may be explained by STS22’s higher word count (461 vs. 61 for other STS datasets; see Table 1). Qualitative observations suggest that the longer texts in STS22 are more robust to lexical changes, whereas paraphrases of shorter sentences rely heavily on synonym replacement. In these shorter contexts, individual tokens carry disproportionate weight, potentially distorting the embeddings. For RubQ we believe this is partly due to a higher robustness of the reranking task as also observed for the English datasets (Table 9, Table 10). Moreover, RubQ’s candidate documents have a substantially higher average word count (63 words vs. 9 for AskUbuntuDuplicateQuestions).
@@
Furthermore, English datasets with initial scores $\geq 80$ (e.g. BIOSSES, SICK-R, STS13/15/17/B, and Banking77) showed relatively large performance gaps (many models -3 to -6pts; also see Figure 7 in the Appendix). This suggests that the initially high scores may depend on shortcuts in token space which are neutralised by paraphrasing. It shows that PTEB’s new evaluation method adds a stress test for semantic invariance, mitigating concerns that MTEB may favour lexical shortcuts and overfitted embedding models. However, neither do we claim that any specific model is more overfitted than another one nor do we make general claims on the performance of individual models as the aim of our study is to present a novel evaluation method; not a comprehensive benchmarking study.
@@
Unexpectedly, smaller encoders were not more sensitive to paraphrasing to a certain extend: Our ablation study on the Qwen3 encoder size showed the similar performance for the 4B vs. 8B parameter models (Table 8). For the datasets in our study, this indicates that robust performance under paraphrase stress can be achieved with smaller embedding models. Furthermore, the prompt sensitivity analysis showed that PTEB is robust to prompt changes further mitigating potential concerns that the results might be brittle due to the use of generative LLMs (Table 12).